Find all solutions in the interval [0, 2π).
cos^2(x) + 2 cos x + 1 = 0
Find all solutions in the interval [0, 2π).
(sin x)(cos x) = 0
Find all solutions to the equation.
cos^2(x) + 2 cos x + 1 = 0

Answer by Alan3354(69443)

Find all solutions in the interval [0, 2π).
cos^2(x) + 2 cos x + 1 = 0
(cos(x) + 1)^2 = 0
cos(x) = -1
x = pi radians
=========================
Find all solutions in the interval [0, 2π).
(sin x)(cos x) = 0
sin(x) = 0
x = 0, pi
------
cos(x) = 0
x = pi/2, 3pi/2